# MORPH: PDE Foundation Models with Arbitrary Data Modality

### v3

@@ Abstract
We introduce MORPH, a modality-agnostic, autoregressive foundation model for partial differential equations (PDEs). MORPH is built on a convolutional vision transformer backbone that seamlessly handles heterogeneous spatiotemporal datasets of varying data modality (1D–3D) at different resolutions, and multiple fields with mixed scalar and vector components. The architecture combines (i) component-wise convolution, which jointly processes scalar and vector channels to capture local interactions, (ii) inter-field cross-attention, which models and selectively propagates information between different physical fields, (iii) axial attentions, which factorize full spatiotemporal self-attention along individual spatial and temporal axes to reduce computational burden while retaining expressivity. We pretrain multiple model variants on a diverse collection of heterogeneous PDE datasets and evaluate transfer to a range of downstream prediction tasks. Using both full-model fine-tuning and parameter-efficient low-rank adapters (LoRA), MORPH outperforms models trained from scratch. Across extensive evaluations, MORPH matches or surpasses strong baselines and recent state-of-the-art models. Collectively, these capabilities present a flexible and powerful backbone for learning from the heterogeneous and multimodal nature of scientific observations, charting a path toward scalable and data-efficient scientific machine learning. The source code, datasets, and models are publicly available at https://github.com/lanl/MORPH.
@@ 1 Introduction
Many problems in physics are governed by partial differential equations (PDEs) where field variables exhibit complex spatiotemporal evolution. Machine learning-based surrogate models have made substantial progress in learning such spatiotemporal physical systems. Prominent standalone surrogates include neural operators such as DeepONet (lu2021learning; kontolati2024learning), Fourier Neural Operators (li2020fourier), the U-shaped Neural Operator (rahman2022u) and transformer-based models (solera2024beta; li2022transformer). Inspired by the success of foundation models in natural language processing, recent research directions in scientific machine learning are transitioning from task-specific to task-agnostic modeling. We use the term “PDE foundation model” to denote a universal PDE surrogate or a universal PDE operator pretrained on large, diverse datasets spanning multiple physics. With task-specific fine-tuning, such models can outperform standalone surrogates and demonstrate utility in data- and compute-scarce scenarios.
@@
Partial observability of physics is a bottleneck in learning a PDE foundation model. Two forces collide in practice. Physics: dynamical systems couple nonlinear, multiscale interactions across scalar and vector fields evolving in four-dimensional continuous space-time, often with widely separated characteristic scales. Data: unlike web-scale text or images, pretraining corpora for PDEs are scarce, costly to curate, and frequently provide only partial observability. The existing benchmark spatiotemporal PDE datasets (takamoto2022pdebench; ohana2024well; herde2024poseidon) contain varying data modality (1D–3D) including trajectories, fields, components and resolutions curated through computationally costly simulations. The datasets are generated under restricted parameter settings with limited variation in initial and boundary conditions. In addition, such datasets can be terabytes in size, making naive homogenization or padding prohibitive. On the experimental side, measuring such fields demands substantial investment and dedicated facilities (e.g., wind tunnels for fluid-dynamics experiments). Most measurement modalities also provide partial observability: seismometers, for instance, yield 1D time series on sparse spatial grids, whereas seismic waves propagate as a (3+1)(3{+}1)-D disturbance through Earth’s crust. Similarly, pressure probes and velocimetry typically provide 1D or 2D temporal observations at limited spatial locations, even though the underlying fluid physics exhibits full spatiotemporal fluctuations in the medium. These practical constraints of simulations and measurements motivate a design that can learn effectively from partial information and generalize across data heterogeneity.
@@
A modeling challenge under partial observations. A foundation model with a transformer backbone provides flexibility in learning from partial physical information. Recent PDE foundation models have made substantial progress toward a task-agnostic universal surrogate. However, most of the works (mccabe2024multiple; chen2024data; herde2024poseidon) assume largely homogeneous data modality, with predominantly 2D Cartesian grids, fixed scalar or vector components of field variables, and limited multiphysics coverage. The architectural choices are not designed for data heterogeneity and scale unfavorably as data modality changes and resolution increases often causing memory/compute blow-ups while requiring task-specific reconfiguration during fine-tuning. In particular, their pretraining sets exclude 3D datasets because patching volumetric data expands the token sequence length, driving quadratic self-attention cost $\mathcal{O}(L_{\text{seq}}^{2})$. The same scaling bottleneck discourages inclusion of higher-resolution 2D datasets. Conversely, 1D problems are often excluded because 2D-centric pipelines enforce padding or tiling up to a 2D layout, wasting memory and compute without adding information. Since many experimental measurements are practically 1D (e.g., point probes, line scans, seismometer traces) and provide the most accurate observation of the underlying physics, we regard 1D as a crucial modality. We therefore treat heterogeneity and multimodality as first-class design requirements for a PDE foundation model. Concurrent works like (ye2024pdeformer; mccabe2024multiple) explicitly highlight the open challenge of supporting more complex domains and extending beyond 2D, underscoring a persistent heterogeneity gap. We identify this as a key bottleneck in current PDE foundation model research. In short, there is a growing need for a flexible and scalable foundation model that can learn from diverse forms of partially observed scientific data.
@@ 2 Related Work
A wide range of ML-based standalone surrogate models are deployed for solving PDEs. Some of these methods are 3D-CNN (wandel2021teaching), ConvLSTM (shi2015convolutional), DCGAN (cheng2020data) and GNNs (chen2021graph). Beyond purely data-driven surrogates, physics-informed neural networks (PINNs) have been applied to spatiotemporal problems (RAISSI2019686). Recently, neural operators like DeepONet (goswami2022deep), Fourier Neural Operators (FNOs) (li2020fourier), wavelet neural operators (WNOs) (tripura2023wavelet), Laplace neural operators (cao2024laplace) and physics-informed neural operators (PINOs) (li2024physics) have gained attention for solving spatiotemporal problems. Another line of methods uses generative models to capture spatiotemporal data distributions. Generative Adversarial Networks (GANs) have been explored to reconstruct or super-resolve turbulent flow fields from partial observations (buzzicotti2021reconstruction). More recently, diffusion models with spatial cross-attention (score-based generative models) have gained traction due to their stability and high fidelity in representing complex distributions (zhuang2025spatially; li2024learning; gao2025generative). Researchers have investigated latent evolution models where spatial learning is performed through autoencoders (oommen2022learning; vlachas2022multiscale; maulik2021reduced; kontolati2024learning) and variational autoencoders (solera2024beta; rautela2024conditional). Temporal dynamics in the latent space have been modeled with recurrent neural networks including LSTMs (maulik2021reduced; vlachas2022multiscale; liu2022hierarchical; rautela2024conditional), Koopman with non-linear forcing (khodkar2021data), DeepONets (oommen2022learning; kontolati2024learning), ODE-Nets (chen2018neural) and transformers with self-attention and easy-attention (solera2024beta; alkin2024universal).
@@
Standalone PDE surrogates are typically specialized to one equation or parameterized family, which means they must be retrained from scratch whenever the PDE form or conditions change. This lack of generality leads to significant computational overhead without knowledge transfer across PDEs. Some initial works have successfully demonstrated the capability of transfer learning for PDEs (goswami2022deep; xu2023transfer). Foundation models for PDEs have emerged to overcome these limitations by capturing the common patterns underlying many different PDE systems. A PDE foundation model is a task-agnostic generalist model pretrained on diverse data, followed with minimal task-specific finetuning. Recent works (yang2023context; chen2024data; herde2024poseidon; chen2018neural; liu2024prose; ye2024pdeformer) demonstrated that a PDE foundation model can outperform standalone models. The advantages in generality, data efficiency, transferability, and versatility motivate the development of foundation PDE models over traditional one-off surrogates (subramanian2023towards). Despite rapid progress, building general-purpose PDE foundation models remains a formidable challenge. PDEs capture complex physics and the datasets representing such spatiotemporal behavior involve several human experts and many compute hours. Public benchmarks remain scarce and frequently provide only partial observability of the underlying physics (takamoto2022pdebench; ohana2024well; herde2024poseidon). It becomes absolutely necessary to build a PDE foundation model with data heterogeneity and modalities in mind (ye2024pdeformer; mccabe2024multiple; morel2025disco). The architectural design should support learning from partial experimental/simulation data and accommodate 1D–3D domains, low to high resolutions, and mixed scalar and vector fields in single and multiphysics settings. However, most of the current PDE foundation models lack this ability.
@@ 3.1 Unified Physics Tensor Format (UPTF-7)
The most direct way to handle data heterogeneity is to pad the datasets to a common shape. However, raw PDE datasets can occupy gigabytes to terabytes of storage. For instance, The Well alone contains $\sim$15 TB of physics simulations (ohana2024well). This makes naive padding both compute and storage-wise inefficient. We therefore seek a unified dataset format that generalizes across PDE datasets while preserving the semantics of the physical observations. We observe that popular benchmarks like PDEBench (takamoto2022pdebench), The Well (ohana2024well), PDEGym (herde2024poseidon), PDEArena (gupta2022towards), and CFDBench (CFDBench) can be represented with a generic 7D data shape. We denote it as $(N,T,F,C,D,H,W)$, where $N$ denotes the number of trajectories, $T$ are the number of time steps, $N\times T$ are the number of snapshots, $F$ are the number of field variables, $C$ are the associated scalar/vector components per field, and $(D,H,W)$ are the spatial depth, height, and width. We retain each benchmark’s native on-disk format and convert mini-batches on the fly to our Unified Physics Tensor Format (UPTF-7) when loading to GPUs. For instance, the 2D compressible CFD data in PDEBench contains velocity, pressure, and density fields where velocity is vector-valued while the other two are scalar. Under our convention, scalars are special cases of vectors. Therefore, we broadcast each scalar component to the maximum vector components count when a batch contains mixed scalar–vector fields. Therefore, the 2D compressible CFD batch maps to a UPTF-7 layout (B,T,F=3,C=2,D=1,H,W)(B,T,F{=}3,C{=}2,D{=}1,H,W). Similarly, a 1D Diffusion–Reaction batch with native $(B,T,W)$ layout maps to UPTF-7 as (B,T,F=1,C=1,D=1,H=1,W)(B,T,F{=}1,C{=}1,D{=}1,H{=}1,W), and a 3D MHD batch in The Well maps to (B,T,F=3,C=3,D,H,W)(B,T,F{=}3,C{=}3,D,H,W). We defer more dataset-specific details to Appendix A.4.
@@
The convolutional operator acts on the component channel(s), producing a set of feature maps determined by the number of filters. Adding convolutional inductive bias helps vision transformers capture locality and improves sample efficiency (dosovitskiy2020image). We can scale feature maps and ultimately MORPH via the number of convolutional filters (default: 8). We partition feature maps into patches of size 8. Across pretraining and fine-tuning datasets, we use at most 4,096 patches, yielding a 32,768-dimensional feature vector, which we linearly project to an embedding whose dimensionality ranges from 256 to 1,024 depending on the model variant.
In most PDE datasets, multiple fields with different spatial scales evolve at various temporal rates. Field-wise multi-head cross-attention prioritizes relevant inter-field interactions and selectively propagates information, enabling careful feature fusion while suppressing spurious activations arising from scale mismatches among coupled fields. The cross-attention operator outputs a single fused field, significantly reducing computational cost (nguyen2023climax). We scale MORPH further via the cross-attention dimension, which by default matches those of axial self-attention.
We factorize full 4D spatiotemporal attention for a (3+1)D dynamical system into time-wise, depth-wise, height-wise, and width-wise axial attentions. For high-dimensional datasets, this limits the computational cost to $O(T^{2}+D^{2}+H^{2}+W^{2})$, in contrast to full spatiotemporal attention with cost $O((TDHW)^{2})$ (ho2019axial). Time-wise axial attention supports flexible temporal context allowing fixed or variable order autoregression (default: 1 step). More details on modeling aspects are deferred to Appendix B.1.
@@ 3.3 Transfer across data modalities
[…]
Here, $E_{\mathrm{PT}}$ is the zero-shot error of the CFD2D-IC pretrained model evaluated on a given target, $E_{\mathrm{SC}}$ is the error of an identical architecture trained fully from scratch on that target, and $E_{\mathrm{NB}}$ is the error of a training-free naive baseline (the same architecture with random initialization and no learning, informally corresponding to near–coin-toss performance), taken as the minimum over 25 random seeds. We define $1-E_{\mathrm{PT}}/E_{\mathrm{NB}}$ as the naive-baseline gain of the pretrained model (NBG-PT), and $1-E_{\mathrm{SC}}/E_{\mathrm{NB}}$ as the naive-baseline gain of the trained-from-scratch model (NBG-SC). In simpler terms, GCR is the ratio $\mathrm{NBG\text{-}PT}$ and $\mathrm{NBG\text{-}SC}$. Our definition of zero-shot GCR is inspired by performance-gap-recovered (PGR) metric introduced in burns2023weak.

### v4

@@ Abstract
We introduce MORPH, a modality-agnostic, autoregressive foundation model for partial differential equations (PDEs). MORPH is built on a convolutional vision transformer backbone that seamlessly handles heterogeneous spatiotemporal datasets of varying data modality (1D–3D) at different resolutions, and multiple fields with mixed scalar and vector components. The architecture combines (i) component-wise convolution, which jointly processes scalar and vector channels to capture local interactions, (ii) inter-field cross-attention, which models and selectively propagates information between different physical fields, (iii) axial attentions, which factorize full spatiotemporal self-attention along individual spatial and temporal axes to reduce computational burden while retaining expressivity. We pretrain multiple model variants on a diverse collection of heterogeneous PDE datasets and evaluate transfer to a range of downstream prediction tasks. Using both full-model fine-tuning and parameter-efficient low-rank adapters, MORPH outperforms models trained from scratch. Across extensive evaluations, MORPH matches or surpasses strong baselines and recent state-of-the-art models. Collectively, these capabilities present a flexible and powerful backbone for learning from the heterogeneous and multimodal nature of scientific observations, charting a path toward scalable and data-efficient scientific machine learning. The source code, datasets, and models are publicly available at https://github.com/lanl/MORPH.
Machine Learning, ICML
@@ 1 Introduction
Many problems in physics are governed by partial differential equations (PDEs) where field variables exhibit complex spatiotemporal evolution. Machine learning-based surrogate models have made substantial progress in learning such spatiotemporal physical systems. Prominent standalone surrogates include neural operators such as DeepONet (Lu et al., 2021; Kontolati et al., 2024), Fourier Neural Operators (Li et al., 2020), the U-shaped Neural Operator (Rahman et al., 2022) and transformer-based models (Solera-Rico et al., 2024; Li et al., 2022). Inspired by the success of foundation models in natural language processing, recent research directions in scientific machine learning are transitioning from task-specific to task-agnostic modeling. We use the term “PDE foundation model” to denote a universal PDE surrogate or a universal PDE operator pretrained on large, diverse datasets spanning multiple physics. With task-specific fine-tuning, such models can outperform standalone surrogates and demonstrate utility in data- and compute-scarce scenarios.
@@
Data heterogeneity is a bottleneck in learning a PDE foundation model. Two forces collide in practice. Physics: dynamical systems couple nonlinear, multiscale interactions across scalar and vector fields evolving in four-dimensional continuous space-time, often with widely separated characteristic scales. Data: unlike web-scale text or images, pretraining corpora for PDEs are scarce, costly to curate, and frequently provide multiple modalities. The existing benchmark spatiotemporal PDE datasets (Takamoto et al., 2022; Ohana et al., 2024; Herde et al., 2024) contain varying data modality (1D–3D) including trajectories, fields, components and resolutions curated through computationally costly simulations. In addition, such datasets can be terabytes in size, making naive homogenization or padding prohibitive. Experiments typically provide sparse, time series or low-dimensional slices of high-dimensional dynamics. For example, seismometers record 1D time series at sparse locations while seismic waves propagate as a (3+1)(3{+}1)-D field; similarly, pressure probes and velocimetry provide 1D/2D slices of fully spatiotemporal fluid dynamics. These constraints motivate models that learn from incomplete measurements and generalize across heterogeneous data or modalities.
@@
A modeling challenge under multiple modalities. A foundation model with a transformer backbone provides flexibility in learning from partial physical information. Recent PDE foundation models have made substantial progress toward a task-agnostic universal PDE surrogate. However, most existing approaches (McCabe et al., 2024; Chen et al., 2024; Herde et al., 2024) implicitly assume near-homogeneous inputs typically 2D Cartesian grids with fixed scalar/vector components and limited multiphysics coverage. As a result, they scale poorly across modalities and resolutions: volumetric patching makes 3D prohibitively expensive by inflating sequence length and the quadratic attention cost $\mathcal{O}(L_{\text{seq}}^{2})$, which also discourages high-resolution 2D pretraining. At the other extreme, 1D data are often forced into 2D pipelines via padding or tiling, wasting compute without adding information. We treat data heterogeneity or multi-modality as first-class requirements. This gap is highlighted by concurrent work (Ye et al., 2024; McCabe et al., 2024) and remains a key bottleneck, motivating a flexible, scalable PDE foundation model that can learn across diverse scientific modalities.
@@ 2 Related Work
A wide range of ML-based standalone surrogate models are deployed for solving PDEs. Some of these methods are 3D-CNN (Wandel et al., 2021), ConvLSTM (Shi et al., 2015), DCGAN (Cheng et al., 2020) and GNNs (Chen et al., 2021). Beyond purely data-driven surrogates, physics-informed neural networks (PINNs) have been applied to spatiotemporal problems (Raissi et al., 2019). Recently, neural operators like DeepONet (Goswami et al., 2022), Fourier Neural Operators (FNOs) (Li et al., 2020), wavelet neural operators (WNOs) (Tripura and Chakraborty, 2023), Laplace neural operators (Cao et al., 2024) and physics-informed neural operators (PINOs) (Li et al., 2024b) have gained attention for solving spatiotemporal problems. Another line of methods uses generative models to capture spatiotemporal data distributions (Buzzicotti et al., 2021; Zhuang et al., 2025; Li et al., 2024a; Gao et al., 2025). Researchers have investigated latent evolution models where spatial learning is performed through autoencoders (Oommen et al., 2022; Vlachas et al., 2022; Maulik et al., 2021; Kontolati et al., 2024) and variational autoencoders (Solera-Rico et al., 2024; Rautela et al., 2024, 2025). Temporal dynamics in the latent space have been modeled with recurrent neural networks including LSTMs (Maulik et al., 2021; Vlachas et al., 2022; Liu et al., 2022; Rautela et al., 2024), Koopman with non-linear forcing (Khodkar and Hassanzadeh, 2021), DeepONets (Oommen et al., 2022; Kontolati et al., 2024), ODE-Nets (Chen et al., 2018) and transformers with self-attention and easy-attention (Solera-Rico et al., 2024; Alkin et al., 2024).
@@
Standalone PDE surrogates are typically specialized to one equation or parameterized family, which means they must be retrained from scratch whenever the PDE form or conditions change. This lack of generality leads to significant computational overhead without knowledge transfer across PDEs. Some initial works have successfully demonstrated the capability of transfer learning for PDEs (Goswami et al., 2022; Xu et al., 2023). Foundation models for PDEs have emerged to overcome these limitations by capturing the common patterns underlying many different PDE systems. A PDE foundation model is a task-agnostic generalist model pretrained on diverse data, followed with minimal task-specific finetuning. Recent works (Yang et al., 2023; Chen et al., 2024; Herde et al., 2024; Chen et al., 2018; Liu et al., 2024; Ye et al., 2024) demonstrated that a PDE foundation model can outperform standalone models. The advantages in generality, data efficiency, transferability, and versatility motivate the development of foundation PDE models over traditional one-off surrogates (Subramanian et al., 2023).
Despite rapid progress, building general-purpose PDE foundation models remains a formidable challenge. PDEs capture complex physics and the datasets representing such spatiotemporal behavior involve several human experts and many compute hours. Public benchmarks remain scarce and frequently provide multiple modalities of the underlying physics (Takamoto et al., 2022; Ohana et al., 2024; Herde et al., 2024). It becomes absolutely necessary to build a PDE foundation model with data heterogeneity and modalities in mind (Ye et al., 2024; McCabe et al., 2024; Morel et al., 2025). The architectural design should support learning from partial experimental and simulation data and accommodate 1D–3D domains, low to high resolutions, and mixed scalar and vector fields in single and multiphysics settings. However, most of the current PDE foundation models lack this ability.
@@ 3.1 Unified Physics Tensor Format (UPTF-7)
The most direct way to handle data heterogeneity is to pad the datasets to a common shape. However, raw PDE datasets can occupy gigabytes to terabytes of storage. For instance, The Well alone contains $\sim$15 TB of physics simulations (Ohana et al., 2024). This makes naive padding both compute and storage-wise inefficient. We therefore seek a unified dataset format that generalizes across PDE datasets while preserving the semantics of the physical observations. We observe that popular benchmarks like PDEBench (Takamoto et al., 2022), The Well (Ohana et al., 2024), PDEGym (Herde et al., 2024), PDEArena (Gupta and Brandstetter, 2022), and CFDBench (Yining et al., 2023) can be represented with a generic 7D data shape. We denote it as $(B,T,F,C,D,H,W)$, where $B$ denotes the batch-size, $T$ are the number of time steps, $B\times T$ are the number of snapshots, $F$ are the number of field variables, $C$ are the associated scalar/vector components per field, and $(D,H,W)$ are the spatial depth, height, and width. We retain each benchmark’s native on-disk format and convert mini-batches on the fly to our Unified Physics Tensor Format (UPTF-7) when loading to GPUs. For instance, batch from 1D-Diffusion–Reaction dataset of PDEBENCH with native $(B,T,W)$ layout maps to UPTF-7 as (B,T,F=1,C=1,D=1,H=1,W)(B,T,F{=}1,C{=}1,D{=}1,H{=}1,W), and a 3D-Magnetohydrodynamics batch of The Well maps to (B,T,F=3,C=3,D,H,W)(B,T,F{=}3,C{=}3,D,H,W). We defer more dataset-specific details to Appendix A.5.
@@
The convolutional operator acts on the component channel(s), producing a set of feature maps determined by the number of filters. Adding convolutional inductive bias helps vision transformers capture locality and improves sample efficiency (Dosovitskiy et al., 2020). We can scale feature maps and ultimately MORPH via the number of convolutional filters (default: 8). We partition feature maps into patches of size 8. Across pretraining and fine-tuning datasets, we use at most 4,096 patches, yielding a 32,768-dimensional feature vector, which we linearly project to an embedding whose dimensionality ranges from 256 to 1,024 depending on the model variant.
In most PDE datasets, multiple fields with different spatial scales evolve at various temporal rates. Field-wise multi-head cross-attention prioritizes relevant inter-field interactions and selectively propagates information, enabling careful feature fusion while suppressing spurious activations arising from scale mismatches among coupled fields. The cross-attention operator outputs a single fused field, significantly reducing computational cost (Nguyen et al., 2023). We scale MORPH further via the cross-attention dimension, which by default matches those of axial self-attention.
We factorize full 4D spatiotemporal attention for a (3+1)D dynamical system into time-wise, depth-wise, height-wise, and width-wise axial attentions. For high-dimensional datasets, this limits the computational cost to $O(T^{2}+D^{2}+H^{2}+W^{2})$, in contrast to full spatiotemporal attention with cost $O((TDHW)^{2})$ (Ho et al., 2019). Time-wise axial attention supports flexible temporal context allowing fixed or variable order autoregression (default: 1 step). We defer further modeling and implementation details to Appendix B.1, and validate the necessity of each design choice (component-wise convolution, field fusion, and axial attention) through the ablations in Sec. 4.4.
@@ 3.3 Transfer across data modalities
$\mathrm{GCR}=\frac{1-(E_{\mathrm{PT}}/E_{\mathrm{NB}})}{1-(E_{\mathrm{SC}}/E_{\mathrm{NB}})}$ (1)
Here, $E_{\mathrm{PT}}$ is the zero-shot error of the CFD2D-IC pretrained model evaluated on a given target, $E_{\mathrm{SC}}$ is the error of an identical architecture trained fully from scratch on that target, and $E_{\mathrm{NB}}$ is the error of a training-free naive baseline (the same architecture with random initialization and no learning, informally corresponding to near–coin-toss performance), taken as the minimum over 25 random seeds. We define $1-E_{\mathrm{PT}}/E_{\mathrm{NB}}$ as the naive-baseline gain of the pretrained model (NBG-PT), and $1-E_{\mathrm{SC}}/E_{\mathrm{NB}}$ as the naive-baseline gain of the trained-from-scratch model (NBG-SC). In simpler terms, GCR is the ratio $\mathrm{NBG\text{-}PT}$ and $\mathrm{NBG\text{-}SC}$. Our definition of zero-shot GCR is inspired by performance-gap-recovered (PGR) metric introduced in (Burns et al., 2023).